One of the most prominent spectral lines of Calcium is the one with

wavelength πœ† = 4226.73 𝐴

π‘œ

[transition from 𝑝(𝑙 = 1) to 𝑠(𝑙 = 0)]. Calcium

atoms exhibit normal Zeeman patterns when placed in a uniform magnetic

field of 4 π‘‡π‘’π‘ π‘™π‘Ž. Calculate the wavelength of the three components of a

normal Zeeman pattern and the separation between them.


1
Expert's answer
2021-12-16T14:39:38-0500

"\lambda=\sqrt{\lambda_x^2+\lambda_y^2+\lambda_z^2},"

"\lambda_x=\lambda_y=\lambda_z,"

"\lambda=\sqrt 3\lambda_x,"

"\lambda_x=\lambda_y=\lambda_z=\frac{\lambda}{\sqrt 3}=2440.30~Å."